OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 5, -3, and -1 + 2i

OpenStudy (anonymous):

sorry it went away

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

before we begin is it clear that two factors are \((x-5)\) and \((x+3)\)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so the real job (before multiplying out) is the find the quadratic polynoimial with zeros at \(-1+2i\) and its conjugate \(-1-2i\)

OpenStudy (anonymous):

you want the easy way, or the real real easy way?

OpenStudy (anonymous):

real real easy

OpenStudy (anonymous):

ok actually lets to the easy way first then the real real easy way

OpenStudy (anonymous):

we can work backwards starting with \[x=-1+2i\] add 1 and get \[x+1=2i\] then square (carefully) to get \[(x+1)^2=(-2i)^2\] or \[x^2+2x+1=-4\]

OpenStudy (anonymous):

add 4 to both sides and get \[x^2+2x+5=0\] and that is your polynomial

OpenStudy (anonymous):

the real real easy way requires memorizing something that if \(a+bi\) is a zero of a quadratic, then it is \[x^2-2ax+(a^2+b^2)\] so in your case \(a=-1,b=2\) and the quadratic is \[x^2-2\times (-1)x+(-1)^2+2^2\] i.e. \[x^2+2x+5\]

OpenStudy (anonymous):

those arent any options tho :(

OpenStudy (anonymous):

your final job is to multiply \[(x-5)(x+3)(x^2+2x+5)\]

OpenStudy (anonymous):

if it was me, i would cheat so as not to screw up the algebra when multiplying

OpenStudy (anonymous):

you know how to do that?

OpenStudy (anonymous):

wait let me try again

OpenStudy (anonymous):

ok i will leave you to it, then show you how to get the answer for sure

OpenStudy (anonymous):

yeah the answer i get does not match up with my choices

OpenStudy (anonymous):

see if that one does

OpenStudy (anonymous):

Oh thank-you! maybe i was multiplying something wrong!

OpenStudy (anonymous):

that is why i said "cheat" is is easy to make a mistake when multiplying all this muck out

OpenStudy (anonymous):

yw